OpenStudy (anonymous):

4t^2(k+9)7+20ts(k+9)7+25s^2(k+9)7 factored : (k+9)^7(2t-5s)^2 now, will u show me the steps to get these?

OpenStudy (anonymous):

You mean this? \[4t^2(k+9)^7+20ts(k+9)^7+25s^2(k+9)^7\]?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Well: \[=(k+9)^7(4t^2+20ts+25s^2) \] Got this part?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

I know the answer but what's the steps?

OpenStudy (anonymous):

Which step you don't understand. From there, you just need to factor 4t^2+20ts+25s^2

OpenStudy (anonymous):

Yeah. I mean steps onhow you factored it

OpenStudy (anonymous):

Are you sure have written the problem correctly?

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

ok. so what's next with (k+9)7(4t2+20ts+25s2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Find two numbers A and B where A*B=4*25=100 A+B=20

OpenStudy (anonymous):

and .. ? that's how you get (k+9)7(4t2+20ts+25s2)?